OpenStudy (anonymous):

a bullet fired into a large block of wood of uniform resistance hits the block with a velocity 48m/s. it penetrates a distance of 60cm and the velocity is then reduced to 24m/s .it penetrates a further distance of ( in cm)?

OpenStudy (wasiqss):

well first calculare the deaacceleration that is 120a=(48)^2-(24)^2

OpenStudy (wasiqss):

assuming uniform deacceleration

OpenStudy (anonymous):

can we make that assumption? does uniform resistance mean uniform decelerating force?

OpenStudy (wasiqss):

actually its not 120a indeed its 1.2 a

OpenStudy (anonymous):

im not saying it doesnt, i dont actually know

OpenStudy (anonymous):

yes it is uniform decceleration

OpenStudy (wasiqss):

1.2a=24^2-48^2

OpenStudy (wasiqss):

shameer i guess now you can do it

OpenStudy (egenriether):

you can find a ratio of kinetic energies, to see how much energy it took to get 60 cm, even though you don't have the mass (it will divide out) then see how much further you can get with the energy you have left using the conversion factor.

OpenStudy (anonymous):

we dont need to find energy it is kinematics problem (motion)

OpenStudy (anonymous):

wasiqss can u give the equation plzz

OpenStudy (wasiqss):

shameer did you find acceleration

OpenStudy (egenriether):

you're right...I have it now, thanks!

OpenStudy (anonymous):

20 cm..i think..

OpenStudy (wasiqss):

after finding acceleration, use the same equation 2as=v^2-u^2

OpenStudy (wasiqss):

where u=0, v=24

OpenStudy (mos1635):

in deaccelaration movement: S=Vo*t-1/2*a*t^2 (1) and Vf=Vo-a*t (2) if you solve (2) to t and repalce that to (1) you will have S=(Vf^2-Vo^2)/2*a

OpenStudy (anonymous):

hey i get a=1400

OpenStudy (anonymous):

can anyone give a solved equation plzzz

OpenStudy (anonymous):

is my ans correct...i used energy conservation principle..

OpenStudy (anonymous):

yes the answer is 20cm i dont know the steps we have to use kinematics equation

OpenStudy (anonymous):

energy conservation is much simler...just equate the ratio of kinetic energy lost between v=[24,48] and[0,24] with the ratio of work done by block to decrease this energies...since KE loss ratio from 48 to 24 is 1/3 therefore the work done ratio must also be 1/3...work done=F.d f=const so ratio of distances is also 1/3... so 60/3 =20 is the ans..

OpenStudy (anonymous):

here KE loss ratio =